### v1

# MEGConformer: Conformer-Based MEG Decoder for Robust Speech and Phoneme Classification

Abstract
We present Conformer-based decoders for the LibriBrain 2025 PNPL competition, targeting two foundational MEG tasks: Speech Detection and Phoneme Classification. Our approach adapts a compact Conformer to raw 306-channel MEG signals, with a lightweight convolutional projection layer and task-specific heads. For Speech Detection, a MEG-oriented SpecAugment provided a first exploration of MEG-specific augmentation. For Phoneme Classification, we used inverse-square-root class weighting and a dynamic grouping loader to handle 100-sample averaged examples. In addition, a simple instance-level normalization proved critical to mitigate distribution shifts on the holdout split. Using the official Standard track splits and F1-macro for model selection, our best systems achieved 88.9% (Speech) and 65.8% (Phoneme) on the leaderboard, surpassing the competition baselines and ranking within the top-10 in both tasks. For further implementation details, the technical documentation, source code, and checkpoints are available at https://github.com/neural2speech/libribrain-experiments.
1 Introduction
In the emerging fields of deep learning and neuroscience, decoding speech from non-invasive brain signals remains one of the central problems in neural signal processing and brain-computer interfaces (BCIs) panachakel2021; dehgan2025; murad2025. The recently released LibriBrain dataset ozdogan2025 and its associated PNPL 2025 Competition landau2025 provide a unique opportunity to address this challenge by enabling large-scale within-subject modeling of magnetoencephalography (MEG) recordings. LibriBrain contains over 50 hours of continuous MEG data from a single participant listening to the complete Sherlock Holmes corpus, accompanied by precise voice activity and phoneme-level alignments. The competition defines two foundational decoding tasks: Speech Detection, which distinguishes between speech and silence, and Phoneme Classification, which predicts the phoneme being perceived from averaged MEG signals. Both tasks include a Standard and an Extended track, where the Standard track restricts participants to using only the official training data. We participated exclusively in the Standard track.
Our approach explores an initial adaptation of state-of-the-art Automatic Speech Recognition (ASR) architectures to non-invasive neural signals. Specifically, we leverage the Conformer model gulati2020, a convolution-augmented Transformer vaswani2017 that combines the global context modeling of self-attention lin2017 with the local feature extraction strength of convolutional networks lecun2015. Conformer-based architectures have been and continue to be the state-of-the-art in the ASR field burchi2021; kim2022; peng2022; kim2023; rekesh2023. We hypothesize that their ability to capture both temporal and spectral dependencies may serve as a good starting point for MEG-based speech tasks, which also exhibit structured spatiotemporal patterns over multiple timescales.
In this work, we evaluate a Conformer model adapted to the LibriBrain competition tasks. The models process raw MEG windows to predict either speech activity or phoneme identity, sharing the same backbone while differing in input and output processing. To ensure fair comparison and reproducibility, all experiments follow the official data splits and evaluation metrics provided by the organizers landau2025.
Our main contributions are fourfold: (i) a unified Conformer architecture jointly optimized for both LibriBrain tasks; (ii) a robust yet straightforward input normalization method that substantially improves holdout generalization; (iii) an effective MEG-specific augmentation; (iv) and a dynamic grouping strategy to classify averaged samples in the Phoneme Task.
This study demonstrates that adapting modern ASR architectures to MEG decoding yields competitive and robust results in both voice activity detection and phoneme recognition, emphasizing the growing convergence between speech processing and neural decoding research.
2 Related work
Decoding speech from non-invasive neural signals has been explored with both EEG and MEG using supervised learning, from linear baselines to deep architectures. EEG studies reconstruct speech features or spectrograms with CNN/Transformer-style models and subject-independent training, showing steady gains with more data accou2023; xu2024. On MEG, prior work demonstrated phrase and word-level decoding and trial-efficiency using wavelet denoising, CNNs, and transfer learning dash2018_1; dash2018_2; dash2019, as well as compact end-to-end networks adapted to sensor time series sarma2024, and transformer-based neural encoding that links linguistic context to MEG responses klimovichgray2023. Other studies have also started to explore phoneme-level decoding in perception and production modalities, comparing traditional and novel model architectures dezuazo2024; dezuazo2025. More recently, LibriBrain established an unprecedented within-subject scale and standardized tasks for speech detection and phoneme classification ozdogan2025; landau2025. Beyond purely supervised setups, contrastive and self-supervised objectives, along with foundation-model guidance, have advanced non-invasive decoding across perception domains, improving retrieval and generation defossez2023; benchetrit2024; banville2025. Recent work on non-invasive brain-to-text systems combines discriminative decoders with language-model rescoring and cross-dataset scaling to produce significant score improvements jayalath2025.
3 Methods
3.1 Model
We use a single Conformer encoder adapted to MEG for both tracks (Speech Detection and Phoneme Classification). The network takes raw sensor time series (306 channels) preprocessed with the official pipeline (bad-channel interpolation, head-position correction, signal-space separation, notch, and band-pass filtering), where windowed signal segments are downsampled to $250\text{\,}\mathrm{H}\mathrm{z}$. For Speech, we use $2.5\text{\,}\mathrm{s}$ windows (625 samples), and for the Phoneme Task, $0.5\text{\,}\mathrm{s}$ windows (125 samples). A lightweight 1D convolutional projection adapts the 306 MEG channels to the Conformer input size (144), followed by a dropout of p=0.1p{=}0.1, a stack of Conformer blocks, and a linear classifier.
Each Conformer block follows the standard macaron layout: feed-forward, multi-head self-attention, depthwise temporal convolution, and a second feed-forward, with residual connections. We keep the design compact with a hidden size of 144. For Speech, based on Conformer Small, we use 16 layers, 4 heads, a hidden layer dimension of FFNs of 576, and the layer’s depthwise convolution layer with a kernel size of 31. For Phonemes, we created a custom-sized Conformer with 7 layers, 12 heads, FFN dimension of 2048, and a kernel size of 127, to better adapt the model to the smaller dataset.
3.2 Speech detection task specifics
Figure 1: Conformer-based model architecture used for Speech Task.
The Speech model is a binary classifier with a single-logit head, binary cross-entropy loss, and loss-level label smoothing of $0.1$ muller2019. We found that simple, speech-style augmentation was sufficient for this task. Therefore, we develop and apply a light MEG-specific variant of SpecAugment park2019, which we introduce as MEGAugment. It includes two operations: (i) time masking, which zeroes two random temporal spans (max width T=180T{=}180 samples at $250\text{\,}\mathrm{H}\mathrm{z}$) per window; and (ii) bandstop masking, which randomly suppresses narrow frequency bands (theta, alpha, beta, gamma, high-gamma mandal2018) using fourth-order infinite impulse response (IIR) notches with probability p=0.4p{=}0.4. Moreover, input windows slide with a 60-sample stride during training to increase diversity via overlapping.
The final layer outputs a probability of speech, which is smoothed during inference by removing speech sequences shorter than 60 samples ($240\text{\,}\mathrm{m}\mathrm{s}$). The block diagram in Figure 1 summarizes the pipeline.
3.3 Phoneme classification task specifics
The Phoneme model uses a linear classification head that outputs 39 logits and is trained with cross-entropy. At inference, we select the class via argmax. To mitigate imbalance in the Phoneme Task, we use class weights $w_{c}$ with the Inverse of Square root of Number of Samples (ISNS) rule following Equation 1, where $n_{c}$ is the count of class $c$ in the training set and $C$ is the total number of classes.
$w_{c}\propto\frac{1}{\sqrt{n_{c}}},\qquad\frac{1}{C}\sum_{c}w_{c}=1,$ (1)
The architecture (detailed in Figure 2) includes an instance-level input normalization layer directly after the raw input. This layer computes per-sample, per-channel normalization across time with no running statistics. We find this choice essential for stable holdout performance: unlike validation and test, the holdout split exhibits different statistical characteristics, indicating distributional shift (see Appendix B for details). Instance normalization essentially removes amplitude and scale drift between windows and sessions. Consequently, it closes most of the holdout gap, even if it is sometimes slightly sub-optimal on the other splits. Because the competition holdout recordings are pre-averaged over 100 samples to improve SNR, we adapt our training setup accordingly. Particularly, we use a 100-sample dynamic grouping loader that reshuffles groups each epoch, allowing the model to see many independent averages of the same class while preserving temporal locality.
Lastly, we ensemble the five best model seeds to smooth predictions and select the final phonemes using majority voting.
Figure 2: Conformer-based model architecture used for Phoneme Task.
4 Experimental setup
We follow the official LibriBrain data partitions for all experiments, using the provided train, validation, test, and holdout splits. In our experience, re-partitioning or rebalancing the data did not yield consistent improvements, so we retain the original configuration throughout.
Training uses the AdamW optimizer loshchilov2018 with a learning rate of $1{\cdot}10^{-4}$, weight decay of $5{\cdot}10^{-2}$, batch size of 256, and early stopping based on validation F1-macro with a patience of 10. We monitor F1-macro on the validation set to select the final checkpoint. These hyperparameters are shared across both tasks.
Evaluation follows the competition protocol, using F1-macro as the primary metric on test and holdout splits.
All model seeds are trained on individual NVIDIA A100 and H100 GPUs, and each model configuration is trained with ten random seeds for better comparison.
5 Results
Our Conformer-based models achieved competitive results in both tasks, as shown in Table 1. In Speech Detection, our best model reached an F1-macro of 88.9%, surpassing the official baseline (68.0%). In Phoneme Classification, our model reached 65.8% F1-macro and again exceeded the competition baseline (60.4%). These results confirm that Conformer architectures, initially developed for ASR, transfer productively to MEG-based decoding and perform competitively with traditional convolutional and transformer models.
Table 1: F1-macro scores in the holdout for both tasks (Standard track). These are preliminary results in the public leaderboard. Our model and scores are marked in bold.
Model Speech Detection $\uparrow$ Phoneme Classification $\uparrow$
Naive Baseline 45.30% 0.47%
Baseline 68.04% 60.39%
Top-10 88.89% 60.96%
MEGConformer 88.90% 65.83%
Top-1 91.66% 73.82%
Beyond leaderboard scores, several design choices contributed to these results. For Speech Detection, extending the input window from $0.5$ to $2.5\text{\,}\mathrm{s}$ gave the largest gain (+10.8%), followed by adopting Conformer over the SEANet baseline (+9.0%) tagliasacchi2020, and reducing the training stride to 60 (+2.8%); MEGAugment had a negligible net effect (+0.01%). In the Phoneme task, dynamic grouping (vs. fixed groups) improved performance by +13.3%, inverse-square-root class weighting outperformed no loss weighting by +7.6%, the custom Conformer added +0.5%, and ensembling of the best seeds contributed an additional +15.4% on the holdout set. Ultimately, instance-level normalization was essential for holdout generalization (over +200%), outperforming batch (+17.8%) and layer normalization (+88.2%). See Appendix A for further details and statistical analysis.
6 Discussion and conclusion
The Conformer architecture proves suitable to decode the spatiotemporal nature of MEG signals, combining convolutional blocks that capture local temporal structure with self-attention for longer-range dependencies. Using very similar models across both tasks, we achieved top-10 leaderboard performance, with minor input preprocessing differences between Speech and Phoneme decoding. Instance-level normalization is indispensable to mitigate distributional shifts between training and holdout splits in the Phoneme Task, substantially improving generalization. Furthermore, a simple SpecAugment-based park2019 augmentation was effective for speech detection, but showed limited benefit for the more complex, imbalanced phoneme classification task. On the other hand, as LibriBrain is single-subject, cross-subject generalization remains an open challenge for future work.
Looking ahead, adapting speech-model architectures for MEG decoding could open the way toward end-to-end speech reconstruction using sequential objectives such as CTC or seq2seq heads, and even enable speech synthesis. Another promising direction is the use of linguistically grounded feature-based representations kwon2025, which may help tackle data imbalance and improve interpretability by decomposing phoneme classification into binary articulatory features (see Appendix C for the challenges found). Finally, our scaling experiments (Appendix D) indicate that while speech decoding performance starts to saturate, phoneme classification continues to improve with additional data, suggesting further gains could be achieved as larger MEG datasets become available.
Acknowledgments and Disclosure of Funding
The Basque Government funds this research (IKUR-IKA-23/18). The authors thank both the DIPC Supercomputing Center and the technological management body of the Basque Government, EJIE, for their technical and human support. This work was conducted as part of the #neural2speech team. Last but not least, special thanks to Ekain Arrieta for inspiring the application of Conformer and ASR-based architectures to MEG decoding.
Appendix Appendix A Ablation details and statistical significance
For the ablation study, we start from our best-performing model and remove each improvement individually while keeping all other settings fixed. Each ablated variant is therefore identical to the full model except for one modification, allowing us to isolate the contribution of that component.
We report per‐variant means and standard deviation over ten seeds and assess paired differences with the Wilcoxon signed-rank test [wilcoxon1945] over the F1-macro scores in the test set. This non-parametric test does not assume normality and evaluates whether the median of paired score differences differs from zero, indicating a genuine effect rather than random variation. We will use a p-value threshold of 0.01 or lower to determine statistical significance.
Speech task ablation
Table 2: Speech task variant test scores over ten seeds before smoothing. Bold marks the best scores.
[…]
Our model 87.59 $\pm$ 0.70 94.72 $\pm$ 0.23 85.72 $\pm$ 1.19 96.15 $\pm$ 0.32 78.64 $\pm$ 1.02
tmax=0.5 78.13 $\pm$ 1.32 90.62 $\pm$ 0.20 75.58 $\pm$ 1.74 90.59 $\pm$ 0.31 65.87 $\pm$ 1.47
SEANet 79.67 $\pm$ 0.50 90.97 $\pm$ 0.16 77.44 $\pm$ 0.59 90.58 $\pm$ 0.34 67.70 $\pm$ 0.61
No stride 85.18 $\pm$ 0.55 93.06 $\pm$ 0.14 83.59 $\pm$ 1.12 94.35 $\pm$ 0.20 75.01 $\pm$ 0.75
No augment 87.58 $\pm$ 0.55 94.20 $\pm$ 0.14 85.82 $\pm$ 1.17 96.08 $\pm$ 0.21 78.53 $\pm$ 0.79
As shown in Table 2, window extension from $0.5$ to $2.5\text{\,}\mathrm{s}$ yields the largest gain over our default (+10.8% relative), and Conformer outperforms SEANet by +9.0%. Both effects are statistically significant: tmax=0.5 vs. ours ($W=0.0$, $p=0.002$), and SEANet vs. Conformer ($W=0.0$, $p=0.002$). Reducing the training stride to 60 also significantly helps (+2.8%; $W=0.0$, $p=0.002$). Removing augmentation has a small effect (+0.01%) and is not significant ($W=13.0$, $p=0.160$). Although MEGAugment improvements are not significant in the final configuration, it did provide a significant gain of +1.8% in earlier model versions, motivating its inclusion ($W=1.0$, $p=0.004$).
SEANet adjustments.
To match $\text{tmax}=$2.5\text{\,}\mathrm{s}$\$ (625 samples at $250\text{\,}\mathrm{H}\mathrm{z}$), we modify only the temporal downsampler, the third conv1d (k=50k{=}50, s=25s{=}25 in the original), by setting its stride to s=160s{=}160. This keeps the intermediate length at $L_{\text{out}}{=}4$, so the subsequent k=4k{=}4 stem still collapses to $1$ (original: $L_{\text{out}}=\lfloor(125-50)/25\rfloor+1=4$; adjusted: $\lfloor(625-50)/160\rfloor+1=4$). We also align the classifier and loss with the binary task by replacing the final $1{\times}1$ conv1d from 512 to 2 with a single-logit 512 to 1 head and using BCE-with-logits plus label smoothing ($0.1$). Beyond architecture, we match the training and evaluation protocol used for our Conformer: sliding-window training with a stride of 60, MEGAugment, AdamW, early stopping on validation F1-macro, best-checkpoint selection, and validation and test scoring with a stride of 1.
Phoneme task ablation
Table 3: Phoneme task test scores over ten seeds before ensembling. Bold marks the best scores.
[…]
Our model 44.29 $\pm$ 2.84 47.75 $\pm$ 2.70 96.67 $\pm$ 0.31
Fixed groups 38.39 $\pm$ 2.83 40.43 $\pm$ 2.77 93.53 $\pm$ 1.66
No weights 40.92 $\pm$ 3.18 43.88 $\pm$ 3.53 95.87 $\pm$ 1.02
Conformer Small 44.09 $\pm$ 4.56 48.49 $\pm$ 4.19 96.33 $\pm$ 0.60
As shown in Table 3, dynamic grouping (vs. fixed groups) yields the largest improvement (+13.3% relative; $W=1.0$, $p=0.004$). Inverse–square‐root class weighting (our default) outperforms non-weighted loss by +7.6% (not significant with $W=10.0$, $p=0.084$). The custom Conformer provides a small lift over Conformer Small (+0.5%; $W=16.0$, $p=0.844$), consistent with similar capacity but with a better fit to the data characteristics. As noted in the main text, holdout ensembling adds +15.4% but is specific to the competition evaluation.
Appendix Appendix B RMS energy analysis across splits in Phoneme Task
To quantify the scale mismatch between partitions, we computed the per-window root-mean-square (RMS) energy across channels and time as in Equation 2, where $x$ is the MEG window (C=306C{=}306 channels, T=125T{=}125 samples). For validation and test, windows were read with the standard competition preprocessing and standardization enabled. The holdout was read as stored (already standardized on disk). We then formed outline-only histograms using the same bin edges and density normalization across splits.
$\mathrm{RMS}(x)=\sqrt{{mean}\big(x^{2}\big)}\quad\text{for }x\in\mathbb{R}^{C\times T},$ (2)
Figure 3 shows that the holdout distribution is bimodal (two peaks near $0.08$ and $0.62$), with markedly larger dispersion (mean $0.64$, standard deviation $0.22$), whereas validation and test are unimodal and tighter (validation $0.54\pm 0.09$, test $0.51\pm 0.09$). This shift motivated the instance-level normalization used in the Phoneme model to remove amplitude and scale differences window-by-window, thereby improving robustness on the holdout.
Figure 3: RMS energy distribution per split in Phoneme Task.
Appendix Appendix C Phonetic feature decoding challenges
To explore linguistically grounded representations for MEG decoding, we trained separate binary classifiers targeting core phonetic features related to the manner of articulation and voicing without sample averaging. Each model was adapted from the Phoneme Classification configuration described in Section 3.3, using the same Conformer backbone and training hyperparameters, but modified for binary output. Notably, we replaced the 39-class head with a single-logit output, used a binary cross-entropy loss with label smoothing and automatic positive-class weighting following Equation 1 to address class imbalance (based on the Speech Detection model in Section 3.2).
This approach follows the ideas proposed by the PNPL competition authors in their blog on language-inspired phoneme classification [kwon2025], where phonetic features are suggested to handle dataset imbalance and improve infrequent phoneme decoding.
Based on [mesgarani2014], we defined six phonetic features based on manner of articulation: Plosive, Fricative, Affricate, Nasal, Liquid, Glide, and a Voicing feature that distinguishes voiced from voiceless phonemes, including vowels and consonants. Table 4 reports preliminary decoding results using F1-macro, positive-class F1, balanced accuracy, and AUROC (mean $\pm$ std across ten seeds).
Following Appendix A, statistical significance is evaluated with a one-sample Wilcoxon signed-rank test against chance ($0.5$ for F1-macro).
Table 4: Decoding performance for binary phonetic features in the non-averaged test set (in %).
[…]
Speech 76.76 88.10 $\pm$ 0.52 94.42 $\pm$ 0.17 86.38 $\pm$ 0.97 96.38 $\pm$ 0.18
Phoneme 5.37 $\pm$ 0.21 4.00 $\pm$ 1.08 5.86 $\pm$ 0.22 64.35 $\pm$ 1.05
Voicing 77.38 57.77 $\pm$ 0.56 83.23 $\pm$ 1.30 57.38 $\pm$ 0.69 63.88 $\pm$ 0.45
Plosive 18.82 55.58 $\pm$ 0.53 25.54 $\pm$ 1.86 55.31 $\pm$ 0.57 62.09 $\pm$ 1.22
Fricative 18.55 53.78 $\pm$ 0.78 23.85 $\pm$ 2.07 53.70 $\pm$ 0.71 57.71 $\pm$ 1.34
Affricate 0.97 49.69 $\pm$ 0.03 0.00 $\pm$ 0.00 49.98 $\pm$ 0.06 60.29 $\pm$ 1.92
Nasal 11.26 51.74 $\pm$ 0.54 12.98 $\pm$ 1.06 51.67 $\pm$ 0.50 55.97 $\pm$ 1.47
Liquid 7.51 50.80 $\pm$ 1.02 8.16 $\pm$ 3.03 51.02 $\pm$ 0.57 55.81 $\pm$ 1.20
Glide 3.62 52.16 $\pm$ 0.81 6.70 $\pm$ 1.65 51.83 $\pm$ 0.70 60.93 $\pm$ 2.25
Among these features, Voicing showed the strongest decoding signal, approaching 58% F1-macro (significant with $W=55.0$, $p=0.001$). Despite its similarity in label balance to the Speech Task, scores are not on par, implying that decoding some fine-grained speech features may not be straightforward. Manner-based features such as Plosive and Fricative were moderately decodable (both significant with $W=55.0$, $p=0.001$), while Affricate, a composite and low-frequency category (0.97%), remained at chance (not significant with $W=0.0$, $p=1.0$). Attempts to address the affricate issue through ensembling of Plosive and Fricative models, transfer learning from these features, and larger effective batch sizes (via gradient accumulation) did not yield convergence improvements. This supports the view that decomposing phonemes into binary linguistic features alone is insufficient to overcome the data scarcity, as some linguistic features may also be rare enough not to be easily decodable, like affricates here.
Overall, these experiments suggest that feature-based representations are promising for improving interpretability and possibly scaling, but further research is needed to address low-frequency classes and explore multitask formulations that jointly learn shared articulatory subspaces.
Appendix Appendix D Data-size ablation
To examine the effect of dataset scale on decoding performance, we conducted data-size ablations for both tasks using progressively larger subsets of the LibriBrain training data. Figures 4 and 5 show the resulting F1-macro scores as a function of the total hours of MEG data used for training, with shaded bands indicating one standard deviation across random seeds.
Figure 4: Data size ablation for Speech Task.
Figure 5: Data size ablation for Phoneme Task.
For Speech Detection, performance rapidly improved with additional data and then began to saturate, suggesting that the task may already approach its ceiling with a single subject’s data. In contrast, Phoneme Classification, even though with some diminishing returns signs, shows an upward trend with no apparent plateau, implying that larger-scale within-subject recordings could still yield substantial gains. These results align with prior observations [ozdogan2025] that decoding performance continues to scale with the amount of high-quality MEG data available for training.